OpenStudy (anonymous):

Find X and Y value: x=-6+y, -7x+9y=8

OpenStudy (anonymous):

\[x=-6+y\\ -7x+9y=8\] replace \(x\) in the second equation by \)-6+y\) and solve \[-7(-6+y)+9y=8\] for \(y\)

OpenStudy (anonymous):

-6y times 7 plus 9y equals 8 so -42 plus 9y equals 8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

first multiply by distributing the \(-7\) in \[-7(-6+y)+9y=8\]

OpenStudy (anonymous):

42

OpenStudy (anonymous):

42 plus 10y=8

OpenStudy (anonymous):

 You will then get 42-7y+9y=8 Combine like terms -7y+9y=2y So 42+2y=8 Subtract 42 on both sides It will cancel out the 42 on the left 2y=-34 Divide by 2 and you get y=-17